OpenStudy (anonymous):

Given the slope and one point, write the equation of the line in slope intercept form m=3,(-3,5)

OpenStudy (fibonaccichick666):

can you tell me the formula for slope intercept form?

OpenStudy (fibonaccichick666):

and can you tell me the formula for slope-point form?

OpenStudy (anonymous):

i know the answer but can you show me how to get the answer.

OpenStudy (fibonaccichick666):

can you answer my questions first?

OpenStudy (anonymous):

slope intercept form is y=mx+b

OpenStudy (fibonaccichick666):

and slope point?

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (fibonaccichick666):

now, can you plug your info into point slope form?

OpenStudy (anonymous):

everytime i do i get confused. but thats because i dont know where to plug in what.

OpenStudy (fibonaccichick666):

ok, so, what is m=?

OpenStudy (anonymous):

slope?

OpenStudy (fibonaccichick666):

and our slope =?

OpenStudy (anonymous):

y=3x+14

OpenStudy (fibonaccichick666):

no, that's not what I'm asking. I'm asking what the value of the slope is

OpenStudy (anonymous):

i don't know. (-3,5)

OpenStudy (fibonaccichick666):

You don't know the slope? it's given. what does m=?

OpenStudy (fibonaccichick666):

Given the slope and one point, write the equation of the line in slope intercept form m=3 (-3,5)

OpenStudy (anonymous):

i don't know what happened there

OpenStudy (fibonaccichick666):

... I really don't know where you are confused. Can you please elaborate?

OpenStudy (anonymous):

i don't understand how to plug in the numbers

OpenStudy (fibonaccichick666):

oh, well what is your slope? m=? so just put whatever m equals in instead of m. What is (x1,y1)? that's your point. Just put in their values where they are respectively

OpenStudy (fibonaccichick666):

ie. 3+x=4 x=1 3+1=4